5. A merchant has coffee worth $40 a pound that she wishes to mix with 60 pounds of coffee worth $90 a pound to get a mixture that can be sold for $70 a pound. How many pounds of the $40 coffee should be used?
how do i solve this

Answer by Edwin McCravy(20055)

A merchant has coffee worth $40 a pound that she wishes to mix with 60 pounds of coffee worth $90 a pound to get a mixture that can be sold for $70 a pound. How many pounds of the $40 coffee should be used?
how do i solve this
Let x = the number of pounds of cheaper coffee.
Since the merchant is mixing this with more expensive coffee, then
x + 60 = the number of pounds of medium priced mixed coffee.

The equation is gotten from this

 +  =  

(x)($40)  + (60)($90)   = (x + 60)($70) or 


             40x + 5400 = 70(x + 60)


Solve that equation and get x = 40

Answer: 40 pounds of $40 coffee.

Edwin